Consider the following time series data.

Week 1 2 3 4 5 6
Value 19 14 17 12 17 14

Develop a three-week moving average for this time series. Compute MSE and a forecast for week 7. Round your answers to two decimal places.

Week Time Series
Value
Forecast
1 19
2 14
3 17
4 12
5 17
6 14


MSE:

The forecast for week 7:

Use = 0.2 to compute the exponential smoothing values for the time series. Compute MSE and a forecast for week 7. Round your answers to two decimal places.

Week Time Series
Value
Forecast
1 19
2 14
3 17
4 12
5 17
6 14


MSE:

The forecast for week 7:

Compare the three-week moving average forecast with the exponential smoothing forecast using = 0.2. Which appears to provide the better forecast based on MSE?



Explain.

  1. Use trial and error to find a value of the exponential smoothing coefficient that results in a smaller MSE than what you calculated for = 0.2. Find a value of for the smallest MSE. Round your answer to three decimal places.

    alpha=
Answer #1

ANSWER

Please refer below table which presents forecast and MSE basis 3 week moving average :

Week

Value

Forecast

Squared error

1

19

2

14

3

17

4

12

16.67

21.78

5

17

14.33

7.11

6

14

15.33

1.78

7

14.33

SUM =

30.67

It may be noted that :

a) Forecast for period t = ( Value for period t-1 + Value for period t-2 + Value for period t-2) /3

Accordingly forecast for week 7 = 14.33

b) Squared error for period t = ( Forecast for period t – Value for period t ) ^2

Sum of squared error = 30.67

Mean square error ( MSE ) = Sum of squared error / Number of data ( i.e. 3 ) = 30.67/3 = 10.22

Please refer below table for forecast and MSE basis exponential smoothing method

Week

Value   

Forecast   

Squared error

1

19

19.00

0.00

2

14

19.00

25.00

3

17

18.00

1.00

4

12

17.80

33.64

5

17

16.64

0.13

6

14

16.71

7.35

7

16.17

SUM =

67.12

It may be noted :

  1. Ft = alpha x Vt-1   + ( 1 – alpha ) x Ft-1 = 0.2 x Vt-1 + 0.8 x Ft-1

Where, Ft , Ft-1 = Forecasts for period t and t-1 respectively

Vt-1 = Value for period t-1

Alpha = exponential smoothing constant = 0.2

We are also making assumption that forecast for week 1 = 19

Accordingly forecast for week 7 = 16.17

  1. Squared error for period t = ( Forecast for period t – Value for period t ) ^2

Sum of squared error = 67.12

Mean square error ( MSE ) = sum of square error / Number of data ( 6 ) = 67.12/6 = 11.186

==================================

if you have any query please ask me in comment box i am here to helps you don't give direct Thumbs down.if you satisfied my work give Thumbs UP

                                                *****THANK YOU******